A spring of constant K is cut into two parts by ratio 1:2. The spring constant K of larger one is? 
answer is 3K/2 but how?

Grade:12th pass

1 Answers

Khimraj
5 years ago
K / K’ = (2L / 3) / L
K / K’ = 2 / 3
K’ = 3K / 2
K’ = 1.5 K

Force constant of larger piece is 1.5 K
Hope it clears..........................